### v1

# Adaptive Focus Memory for Language Models

Abstract.
Large language models (LLMs) are increasingly deployed in multi-turn dialogue settings, but their behavior is still bottlenecked by fixed context windows and naïve memory strategies. Replaying the full conversation at every turn is simple but expensive, while static summarization or recency-only heuristics often erase safety-critical user details. We present Adaptive Focus Memory (AFM), a dynamic context manager that assigns each past message one of three fidelity levels—FULL, COMPRESSED, or PLACEHOLDER—based on semantic similarity to the current query, half-life recency weighting, and importance classification. AFM packs messages chronologically under a strict token budget, preferring high fidelity for the most relevant turns while aiming to preserve a cheap trace of the dialogue. In a safety-oriented benchmark involving a user with a severe peanut allergy planning a trip to Thailand, AFM retains the allergy across both short and medium-length conversations, matches the safety performance of naive replay, and cuts average token usage by 66% relative to a replay baseline. We release a modular Python implementation of AFM designed for OpenAI-compatible APIs and offline operation, enabling practitioners to reduce inference cost without sacrificing safety or factual continuity in the evaluated scenario.
1. Introduction
Large language models (LLMs) such as ChatGPT, Claude, and Gemini have become powerful general-purpose dialogue agents. However, their performance remains constrained by limited context windows and the high inference-time cost of each token. As conversations grow in length, naïvely replaying the full history at every turn leads to bloated prompts, rising latency, and mounting API expenses—posing serious challenges for multi-agent systems, assistants, and long-form interactions.
To mitigate these inefficiencies, prior work has explored memory-management strategies like Retrieval-Augmented Generation (RAG) and summarization. RAG stores prior content in external vector databases and retrieves relevant chunks per query. Summarization periodically compresses old messages into short abstractions. However, both approaches face limitations in dialogue settings: RAG pipelines require dedicated retrieval infrastructure and often disrupt conversational flow, while static summaries introduce irreversible information loss and cannot adapt to shifting user intent.
Adaptive Focus Memory (AFM) introduces a dynamic alternative. Rather than replaying all history or summarizing indiscriminately, AFM allocates each past message to one of three fidelity levels: FULL (included verbatim), COMPRESSED (summarized via an LLM or heuristic), or PLACEHOLDER (replaced with a short reference stub). This selective replay preserves salient information while minimizing prompt length.
AFM scores messages using a combination of embedding-based semantic similarity, half-life recency decay, and an importance classifier. It can operate in fully offline (heuristic) or LLM-assisted modes, and integrates seamlessly into OpenAI-compatible chat pipelines. In benchmark evaluations on a safety-critical synthetic dialogue scenario, AFM reduces token usage by roughly two-thirds while preserving factual continuity.
Contributions. This work makes the following contributions:
  • We propose a dynamic context selection framework that scores past messages using semantic similarity, temporal recency, and LLM-based importance classification.
  • We introduce a multi-fidelity replay mechanism that assigns each message to FULL, COMPRESSED, or PLACEHOLDER form and packs messages chronologically under a fixed token budget.
  • We demonstrate that AFM retains critical user-specific facts—such as allergy declarations—more reliably than recency-only baselines, while matching the safety performance of naive replay at substantially lower cost.
  • We release an open-source AFM implementationhttps://github.com/cruz209/AFMforLLM suitable for reproducible research and integration into production systems.
The rest of the paper is organized as follows: Section 2 reviews prior work. Section 3 (Method) describes the AFM algorithm. Section 4 (Experiment) presents benchmarks. Section 5 (Discussion) covers limitations and broader implications, and Section 6 concludes.
2. Related Work
2.1. Long-Context Language Models
Early transformer architectures [2] established the self-attention mechanism that enables contextual reasoning but introduced a quadratic memory cost that continues to limit practical context length. Subsequent large-scale generative models such as GPT-3/4 [3, 4] and LLaMA [5] extend context windows but still pay linear inference-time cost proportional to prompt length. Recent industrial systems (e.g., Claude 3 [6]) expand length further, yet running them efficiently remains a bottleneck for multi-turn interaction.
Several works propose architectures designed for local adaptation or long-context specialization. Meng et al. [7] introduce locally trainable memory layers for more efficient fine-tuning, while Geva et al. [12] show that transformer feed-forward layers operate implicitly as key–value memory stores. These approaches improve contextual capacity but do not directly reduce inference cost, which remains proportional to the entire visible context.
2.2. Retrieval-Augmented and External-Memory Approaches
Retrieval-Augmented Generation (RAG) [8] offloads knowledge into external vector stores and selectively injects relevant snippets at inference time. While effective for fact-intensive tasks, RAG often underperforms in conversational settings due to retrieval noise and lack of discourse continuity. Roberts et al. [11] demonstrate the limits of relying purely on internal parameters for knowledge storage, motivating hybrid external-memory architectures.
More recent surveys [9] highlight the difficulty of maintaining coherence when summarizing or retrieving historical dialogue, particularly when the user’s intent evolves across turns. Structured-memory architectures [13] attempt to improve retrieval fidelity, yet these systems require substantial infrastructure—vector databases, indexing, and latency-heavy retrieval pipelines.
2.3. Summarization and Compression of Conversational History
A classical strategy for context control is periodic summarization, either extractive or abstractive. Wu et al. [10] examine long-context summarization to reduce token load, but fixed-frequency summarization introduces irreversible information loss and can cause hallucinations when summaries drift from user intent.
Extractive heuristics and sentence-selection methods remain lightweight but often fail to preserve subtle constraints—especially user-specific preferences or safety-critical details. Abstractive models perform better but at the cost of additional inference steps, which increases cumulative compute.
2.4. Tokenization and Efficient Inference
Inference cost is known to scale roughly linearly with input token count across major LLM stacks. Industry-standard tokenizers such as tiktoken [14] and dataset libraries such as HuggingFace Datasets provide the measurement tools used to benchmark compression efficacy in this work.
Compute-efficient LLM design groups (DeepSeek, OpenAI, Anthropic, Meta) increasingly emphasize token-level cost reduction as a primary engineering constraint. However, most existing work focuses on model-level changes (attention variants, KV caching, or speculative decoding), whereas AFM operates orthogonally: it reduces cost via dynamic, semantic-aware prompt packing without any model modifications.
2.5. Position of AFM Within the Literature
AFM differs from prior approaches in three key ways:
  • Dynamic fidelity assignment: Previous systems often treat all messages uniformly (full retention or global summary), whereas AFM assigns each message one of three fidelity levels—FULL, COMPRESSED, or PLACEHOLDER—based on semantic relevance, recency, and importance.
  • Model-agnostic, plug-and-play design: AFM requires no architectural changes, no retrieval store, and no fine-tuning. It operates entirely at the prompt layer and is implemented as a simple Python library.
  • Explicit compute minimization: Unlike long-context models or pure summarization strategies, AFM explicitly focuses on reducing token-level inference cost and reports these savings in standard units (tokens and dollars per million tokens).
Taken together, AFM occupies a space between RAG, KV-extended transformers, and classical summarization—providing a memory framework whose primary objective is token-efficient factual continuity rather than scaling raw context length or storing external facts.
3. Method
3.1. Overview.
Adaptive Focus Memory (AFM) is implemented as a pluggable context manager that operates entirely at the prompt-construction layer. Rather than replaying the full conversation history at every turn, AFM decides, for each past message, whether to include it verbatim, include a compressed summary, or replace it with a short placeholder stub.
We frame the problem as a heuristic, greedy packing process under a fixed budget . Given a conversation history $H=\{m_{1},\dots,m_{t}\}$, a current query $q_{t}$, and a maximum prompt budget $B$ (in tokens), AFM assigns each message $m_{i}$ a fidelity tier
$f(m_{i})\in\{\textsc{Full},\textsc{Compressed},\textsc{Placeholder}\}$
and then constructs a chronological prompt whose total estimated token length does not exceed $B$, whenever possible. The implementation prioritizes messages that are semantically similar to $q_{t}$, temporally recent, and classified as important by a small LLM-based classifier when an API is available.
3.2. Relevance scoring.
The reference implementation combines three signals in a single scalar score per message: semantic similarity, recency, and an LLM-based importance label.
Semantic similarity.
Each message $m_{i}$ is embedded once (lazily on first use) and cached. Given the embedding $E(m_{i})$ and the embedding $E(q_{t})$ of the current query, AFM computes cosine similarity
$\operatorname{sim}(m_{i},q_{t})=\frac{E(m_{i})\cdot E(q_{t})}{\|E(m_{i})\|\,\|E(q_{t})\|}.$
The reference implementation uses an OpenAI embedding model (text-embedding-3-small) when an API key is configured, and a dependency-free hashing-based embedder otherwise.
Recency weighting.
Let k=tik=t-i be the number of turns since message $m_{i}$ was created and let $h$ be a configurable half-life parameter. AFM computes a recency weight
$w_{\text{recency}}(m_{i})=0.5^{k/h},$
so that the influence of a message decays exponentially as it moves further back in the dialogue. In our experiments we use a fixed half-life $h=12$ turns unless otherwise noted.
Importance classification.
When an OpenAI client is available, AFM calls a small chat model (gpt-4o-mini) once per message to classify its importance as CRITICAL, RELEVANT, or TRIVIAL. The classifier is prompted to treat safety-sensitive facts (e.g., medical conditions) as CRITICAL. If no client is configured, all messages default to TRIVIAL and no classification calls are made. **AFM does not evaluate the accuracy of the importance classifier, its behavior is prompt-driven.**
Final scoring rule.
The actual scoring function in the reference implementation is piecewise, not a single multiplicative expression. Let $\operatorname{sim}_{i}$ be the cosine similarity and $w_{\text{recency},i}$ the recency weight for $m_{i}$. Then:
[…]
Critical messages are thus force-elevated to the maximum score, while relevant and trivial messages use similarity- and recency-weighted scores with different scales. These scores are only used for fidelity decisions; they do not reorder the conversation.
3.3. Fidelity assignment and context packing.
AFM uses the scores $s_{i}$ to assign an intended fidelity tier to each message before packing:
  • If $s_{i}\geq\tau_{\text{high}}$, mark $m_{i}$ as intended Full.
  • Else if $s_{i}\geq\tau_{\text{mid}}$, mark $m_{i}$ as intended Compressed.
  • Else, mark $m_{i}$ as intended Placeholder.
The thresholds $\tau_{\text{high}}$ and $\tau_{\text{mid}}$ are hyperparameters (e.g., $\tau_{\text{high}}=0.45$, $\tau_{\text{mid}}=0.25$ in our OpenAI-backed experiments).
Packing is then performed in strict chronological order. Let $B$ be the token budget and $b_{\text{left}}$ the remaining tokens. For each message $m_{i}$ (from oldest to newest), AFM attempts to include the highest-fidelity representation consistent with its intended tier and the remaining budget:
  • If $m_{i}$ is intended Full, AFM:
    1. (1)
      tries the full text; if the estimated token count fits within $b_{\text{left}}$, it is appended;
    2. (2)
      otherwise, generates or reuses a compressed summary and tries that;
    3. (3)
      if the summary still does not fit, it falls back to a stub.
  • If $m_{i}$ is intended Compressed, AFM attempts the compressed form first and falls back to a stub if the compressed form does not fit.
  • If $m_{i}$ is intended Placeholder, AFM only attempts the stub.
These hyperparameters were chosen heuristically based on a small number of
development conversations; no large-scale tuning was performed.
In all cases, AFM estimates token usage with a shared TokenCounter before committing to adding a representation. If even the stub for a message would exceed the remaining budget, that message is dropped from the prompt. Therefore, AFM attempts to maintain a cheap trace of low-importance turns, but it does not guarantee that every turn is represented under extremely tight budgets.
3.4. Compression layer.
The reference implementation supports two interchangeable compressors via a common Compressor interface:
  • HeuristicCompressor is a fully local, extractive compressor. It operates by ranking sentences using simple lexical overlap with a query hint, mild length penalties, and positional bias. It then truncates to a target token budget based on the shared TokenCounter. This compressor enforces the budget deterministically.
  • LLMCompressor is an abstractive compressor that calls an OpenAI chat model (gpt-4o-mini in our experiments) with a prompt instructing it to rewrite the input under a specified token budget. The compressor relies on the model to follow the budget instruction; the current implementation does not re-check or truncate the returned summary.
AFM selects the compressor at initialization time. In the reference code, OpenAI-backed runs use LLMCompressor; offline runs fall back to HeuristicCompressor.
3.5. Token budgeting and enforcement.
Token counts are estimated by a TokenCounter utility. When the tiktoken library is available, it uses the encoding for gpt-4o-mini and falls back to cl100k_base on error; if tiktoken is not available, it uses a naive whitespace-based word count as a proxy for token length.
During packing, AFM maintains a running budget $b_{\text{left}}$, initialized to the user-provided budget $B$. Every time it proposes to add a representation (full text, compressed summary, or stub), it first estimates the token count with TokenCounter. If the representation fits ($\text{tokens}\leq b_{\text{left}}$), AFM appends it and decrements $b_{\text{left}}$; otherwise, it tries a lower-fidelity form or drops the message entirely if no representation can fit.
A system preamble (if provided) is treated like any other message: it is included first, subject to the same budget accounting.
3.6. System architecture and API.
AFM is implemented as a modular Python package centered on a FocusManager class. The key methods are:
  • add_message(role, content): append a new message to the internal history. Messages are assigned unique IDs, and embeddings are computed lazily.
  • build_context(current_query, budget_tokens, system_preamble): compute relevance scores, assign intended fidelity tiers, and return a list of $(\text{role},\text{content})$ pairs representing the packed prompt, along with summary statistics (e.g., total tokens used, number of compressed messages, etc.).
The same interface works in both online and offline modes and is designed to drop into OpenAI-style chat pipelines with minimal glue code.
3.7. Backend integration.
The reference implementation supports two embedding backends and two compression backends:
  • When an OpenAI API key is available, AFM uses an OpenAIEmbedder that wraps text-embedding-3-small and the LLMCompressor for abstractive compression.
  • When no API key is available, AFM falls back to a local HashingEmbedder and the HeuristicCompressor.
Similarly, importance classification is only performed when an OpenAI client is configured. If the classifier or compressor calls fail due to network or API errors, the current implementation allows the exception to propagate rather than attempting automatic retries or alternative backends.
3.8. Fallback and robustness.
AFM’s robustness features are deliberately minimal in the reference code. The only automatic “fallback” inside the core algorithm is fidelity downgrading: if a message cannot fit in Full form, AFM tries Compressed, then Placeholder. Beyond that, there is no automatic retry logic, rate limit handling, or circuit-breaking. Adding such mechanisms is left for future versions.
3.9. Design philosophy.
AFM is intentionally model-agnostic and infrastructure-light. It requires no vector database, no KV-cache modifications, and no fine-tuning. All state remains in process as a simple list of messages plus per-message metadata (embedding, score, compression state).
In contrast to static summarization approaches that periodically compress the entire history into a single buffer, AFM scores and revisits every message at each call to build_context. This allows it to adapt to changes in user intent over time while still providing explicit, budget-aware control over how much space each past turn occupies in the prompt.
4. Experiments
4.1. Objective and scope.
Our experiments evaluate how effectively AFM heuristically balances three competing goals in multi-turn dialogue: (1) retaining user-specific factual constraints, (2) avoiding unsafe recommendations, and (3) minimizing token usage and latency. We focus on a safety-critical setup where the user has a severe peanut allergy and is planning a trip to Thailand.
4.2. Task design.
We construct a synthetic but realistic conversation where a user with a life-threatening peanut allergy asks about travel logistics, street food, and cultural experiences in Thailand. The allergy is stated early and then not repeated verbatim, while later questions implicitly require the model to remember and apply that constraint (e.g., “the street food sounds AWESOME I wanna have it all” and “What are the best street foods I should try”).
Following the scripts in the benchmark repository, we consider two lengths of this scenario:
  • Short (3 turns). The allergy and food-related queries are close together.
  • Medium (9 turns). The allergy is mentioned early, followed by several intervening questions (e.g., about destinations, transport, and Muay Thai) before the final street-food query.
In both cases, we evaluate whether each method remembers and correctly applies the peanut allergy when recommending food.
4.3. Compared methods.
We compare AFM against three baselines that correspond directly to the provided benchmark scripts:
  1. (1)
    Default (stateless) chat. Only the current user turn and a fixed system prompt are sent to the model. No prior messages are included; there is no token budget enforcement. This is implemented in default.py.
  2. (2)
    Naive truncated replay. All previous user and assistant messages are appended to the prompt, subject to a global token budget $B=800$. When the budget is exceeded, the oldest messages are dropped until the prompt fits. No compression is used; all retained messages are verbatim. This is implemented in naive_replay.py.
  3. (3)
    Recency-based compression. The two most recent user–assistant turns are kept verbatim. All earlier messages are individually compressed to a fixed local budget (100 tokens per message) using LLMCompressor, and the compressed messages plus recent turns are concatenated without an additional global token cap. This is implemented in recency.py.
AFM itself is instantiated via FocusManager with OpenAI embeddings and the LLMCompressor when an API key is present, matching the configuration used by the recency-based baseline.
4.4. Implementation details.
All experiments use OpenAI’s gpt-4o-mini model as the backend for generation and, when applicable, for compression and importance classification. We use the following decoding temperatures:
  • Default (stateless) baseline: temperature $0.0$ (deterministic).
  • Naive truncated replay baseline: temperature $0.7$.
  • Recency-based compression baseline: temperature $0.7$.
  • AFM in our interactive demo: temperature $0.0$ for deterministic compression and context selection.
Token usage in all baselines is measured with the shared TokenCounter configured for gpt-4o-mini. The truncated replay baseline explicitly uses a budget of 800 tokens when deciding which messages to retain; the recency-based baseline and the default baseline do not enforce a global budget and therefore can produce longer prompts.
AFM’s hyperparameters in the OpenAI-backed experiments are:
  • high-fidelity threshold $\tau_{\text{high}}=0.45$,
  • medium-fidelity threshold $\tau_{\text{mid}}=0.25$,
  • recency half-life $h=12$ turns,
  • maximum stub length of 12 tokens,
    These values are drawn directly from the FocusConfig in the reference code.
    4.5. Evaluation metrics.
    We assess each method along the following axes:
    • Allergy retention. Whether the model explicitly recalls and applies the user’s peanut allergy when recommending food (binary pass / fail per conversation).
    • Factual safety. Whether any recommended dishes clearly contain peanuts or common peanut-derived ingredients without appropriate warnings.
    • Token usage. Total tokens consumed per conversation (system prompt, all context, and assistant responses), averaged across the short and medium runs.
    • Latency. Wall-clock time per assistant turn, computed from prompt submission to receipt of the completion.
    Token and latency metrics for the baselines are logged directly by the benchmark scripts. AFM token and latency measurements are collected using the same TokenCounter and timing conventions.
    4.6. Main results.
    Table 1 reports representative results from a single short- and medium-length run per method under the configurations above. The numbers for the three baselines are computed directly from the provided benchmark logs; AFM’s numbers are obtained using the same measurement infrastructure in the AFM demo script.
    Method Allergy Recall (S / M) Avg Tokens Latency (s) Safe?
    Default (stateless) N / N 1,493 6.6 No
    Naive truncated replay Y / Y 2,479 8.8 Yes
    Recency compression Y / N 1,888 12.0 Potentially
    AFM (ours) Y / Y 504 6.2 Yes
    Table 1. Benchmark results on short (S) and medium (M) conversations. Baseline numbers are calculated across the short and medium runs logged by the benchmark scripts; AFM is measured with the same tokenizer and timing setup. Reported metrics come from a single reference run and should be interpreted qualitatively, as AFM relies on external API calls (embedding, compression, importance classification) that make multi-run statistical evaluation costly and noisy due to API and network variance. A more comprehensive, multi-run evaluation is left for future work. Note that the recency-compression baseline does not enforce a global token budget and can exceed the 800-token cap that AFM and naive replay obey. This matches common production deployments of recency heuristics but means the baseline is not a strictly controlled comparison. Latency comparisons are approximate because decoding temperatures differ across baselines.
    Two patterns stand out. First, the stateless and recency-compressed baselines fail to reliably apply the allergy constraint in the medium-length scenario: they eventually drift to generic street-food recommendations that do not reflect the user’s stated constraint. Second, naive truncated replay maintains safety but at the cost of substantially larger prompts and higher latency. AFM achieves the same safety as naive replay on this benchmark while using considerably fewer tokens.
    4.7. Qualitative failure modes.
    Inspecting the transcripts reveals characteristic failure modes:
    • Default (stateless). Because the model only sees the current user message and a generic system prompt, it recommends canonical Thai street foods (including peanut-heavy dishes such as Pad Thai) without referencing the allergy, even when the allergy was stated earlier in the same conversation scenario under other methods.
    • Recency-based compression. When the allergy happens to be within the last few turns, compressed context plus recent raw turns is often sufficient. Once the allergy falls outside the recent window, the model reverts to generic advice such as “if you have food allergies, ask vendors about ingredients,” without explicitly accounting for peanuts.
    • Naive truncated replay. This baseline successfully preserves the allergy but often feeds the model long, repetitive prompts as the history grows. The resulting responses can be verbose and redundant, with little benefit over AFM in the benchmark scenario.
    AFM, by contrast, assigns the allergy declaration a high importance score, securing it a Full slot even when older, and tends to keep the later food-related user turns at higher fidelity as well. This leads to explicit allergy-aware recommendations in both short and medium settings.
    4.8. Ablation analysis (planned).
    The current codebase exposes natural ablation dimensions—removing compression, disabling importance classification, altering the recency half-life, or swapping the LLM compressor for the heuristic one. However, we have not yet run a systematic ablation study. Formal ablations over these dimensions are left to future work and are not reported here.
    4.9. Human evaluation.
    We did not conduct a formal human evaluation with multiple annotators and pre-registered protocols. Instead, we relied on qualitative inspection of the logged transcripts across methods. In these spot checks, AFM’s responses were consistently more explicit about the peanut allergy in the final food recommendation turns, whereas the baselines often produced allergy-agnostic street-food lists once the allergy drifted out of their effective context window.
    Because this inspection was informal and limited in scope, we do not report numerical human preference percentages. A more rigorous human evaluation is future work.
    4.10. Compute cost analysis.
    To relate token savings to monetary cost, we define a simple compute saving ratio (CSR). Let $T_{\text{base}}$ be the average number of tokens consumed by a baseline configuration and $T_{\text{AFM}}$ be the average tokens consumed by AFM on the same conversations. We define
    $\mathrm{CSR}=\frac{T_{\text{base}}-T_{\text{AFM}}}{T_{\text{base}}}.$
    Using the stateless default baseline as $T_{\text{base}}$ (average of short and medium runs), the empirical values in Table 1 yield $\mathrm{CSR}\approx 0.66$, i.e., AFM reduces token usage by roughly 65% relative to a default configuration that does not manage history at all. Relative to the replay-style baselines (naive truncated replay and recency compression), AFM achieves even larger fractional token reductions, on the order of 70–80% in this benchmark.
    Assuming a linear cost model with price $C_{\text{perM}}$ USD per million input tokens—as is standard for commercial LLM APIs—the marginal cost reduction per million tokens is
    $\Delta C=\mathrm{CSR}\times C_{\text{perM}}.$
    For example, with $C_{\text{perM}}=\mathdollar 3$ and $\mathrm{CSR}\approx 0.66$, AFM would save roughly \$2 per million tokens relative to the default baseline. In high-throughput deployments, this cost difference compounds rapidly.
    4.11. Summary.
    On the evaluated allergy benchmark, AFM:
    • preserves safety and allergy recall in both short and medium-length conversations;
    • substantially reduces token usage compared to both a stateless default setup and replay-based baselines;
    • avoids the extreme prompt bloat of naive replay while retaining key user-specific constraints more reliably than a simple recency-only compressor.
    These results support the central claim of AFM: selectively allocating fidelity across past messages can yield strong safety and factual continuity and accuracy at significantly lower token cost than naive history management strategies.
5. Discussion
5.1. When Context Is Not Comprehension
One of the core findings in our experiments is that merely including prior messages in an LLM’s prompt does not guarantee effective memory. The naive replay strategy—common in many production chat systems—preserves all text but still fails to consistently apply user-specific facts when they matter. In our allergy benchmark, stateless and recency-compressed configurations ignored life-critical details (e.g., a peanut allergy) just one or two turns after they appeared.
This reinforces a central problem in long-context inference: information salience degrades with token distance. LLMs, even with extended context windows, exhibit a recency bias and a tendency to overweight immediately proximal text. This creates an illusion of memory where none exists and highlights the need for systems that explicitly curate and scaffold model attention over time.
5.2. AFM’s Role in the Context Stack
AFM intervenes at the level of prompt construction—neither modifying the model weights nor introducing complex retrieval infrastructure. This makes it suitable for developers seeking efficient memory management without sacrificing simplicity, interpretability, or control.
Whereas RAG pipelines introduce database latency and risk discourse in coherence due to chunk fragmentation, AFM operates inline, restructuring the prompt without external calls. Compared to static summarization buffers (e.g., fixed-length history windows or periodic compress-and-truncate), AFM dynamically rescores every message per query, allowing real-time adaptation to user intent.
In this sense, AFM functions as a soft attention mechanism over dialogue history—prioritizing based on both semantic relevance and temporal context, and deciding how “heavy” each memory needs to be: full, compressed, or placeholder. AFM clamps cosine similarity at zero, discarding negative similarity information.
5.3. Efficiency Without Blind Compression
One of AFM’s key contributions is demonstrating that memory compression need not be all-or-nothing. Instead of summarizing the entire past into a single global buffer (which risks flattening important details), AFM preserves structure. Every message is individually scored and passed through a fidelity filter, enabling nuanced context packing.
Empirically, this approach achieves strong token efficiency—reducing prompt size by over 60% on average—without compromising on safety-critical memory. AFM outperforms both naive replay and recency-only baselines on safety recall and token efficiency, and achieves competitive latency in our evaluated scenario.
5.4. Failure Modes and Challenges
AFM, while promising, is not immune to edge cases. Several challenges merit deeper scrutiny:
  • Semantic drift from LLM compression. Abstractive summarization is inherently lossy. While effective for reducing tokens, poor LLM summaries may misrepresent prior user intent or omit subtle qualifiers—especially for long or multi-part messages. Hallucinated details, although rare in our benchmark, remain a risk under low-temperature decoding.
  • Overrepresentation of low-information turns. Messages with high embedding similarity but low informational value (e.g., “sure,” “got it” ) may receive elevated scores unless filtered via length or entropy thresholds. Future versions could incorporate dialogue-act classification to suppress such filler turns.
  • Interaction effects across messages. AFM scores each message independently, ignoring interactions like pronoun chains or deferred clarifications. This may result in dropped context for referential utterances unless earlier turns are retained. A graph-based salience model might help capture such dependencies.
  • Static hyperparameters. Current AFM settings use fixed half-life and importance multipliers. Adaptive weighting (e.g., task-conditioned recency schedules) could better balance short-term versus long-term salience.
5.5. Toward Learned or Reinforced Memory Policies
AFM’s current decision logic is greedy and rule-based. However, as context management becomes more central to multi-agent systems, we envision extensions that learn optimal memory policies:
  • Reinforcement-learned packing. Agents could be rewarded for efficient context usage that preserves factual recall or reduces hallucination rate.
  • Memory trees. Instead of a flat log, context could be stored in a hierarchical structure with collapsible nodes and rewritable summaries.
  • Personalization-aware retention. Messages may be scored differently based on user identity or long-term intent profiles. This would require coupling AFM with identity tracking or goal-prediction modules.
5.6. Broader Implications
As LLMs are increasingly embedded in tool-using agents, IDE copilots, or multi-turn reasoning chains, the need for cost-aware, interpretable, and controllable memory systems is growing. AFM’s design makes it especially attractive in the following real-world scenarios:
  • Conversational agents with safety constraints. Applications like healthcare chatbots or legal assistants must retain critical facts across turns without compromising on latency or privacy.
  • Agentic workflows with recurrent state. Planning agents or goal-oriented bots benefit from a memory that preserves task-relevant state across tool invocations.
  • Model distillation and compression pipelines. AFM can serve as a teacher signal in training smaller models to simulate long-context behavior without full replay.
In all these contexts, the need is the same: not just to remember more, but to remember better. AFM’s selective fidelity framework takes a concrete step toward that goal.
6. Conclusion
Large language models are only as effective as the context they are given, yet the standard practice of naïvely replaying the full dialogue history is increasingly untenable in terms of cost, latency, and memory reliability. Adaptive Focus Memory (AFM) offers a principled alternative: a pluggable modular framework that dynamically determines the fidelity of prior messages based on relevance and recency, while enforcing strict token budgets.
Through a combination of embedding-based scoring, half-life recency weighting, and hybrid compression techniques, AFM preserves the key factual constraint in our evaluated scenario while aggressively reducing prompt size. Our experiments show that AFM outperforms baseline approaches on key metrics of factual retention, safety alignment (AFM matches naive replay on safety alignment in our evaluated scenario at a much lower token cost.), and token efficiency in a safety-critical benchmark. Unlike retrieval-augmented or static summarization methods, AFM operates inline, requires no external infrastructure, and adapts fluidly to changing user intent.
The core insight of AFM—that not all memory deserves equal fidelity—paves the way for more intelligent, interpretable, and efficient use of large models in long-horizon settings. As the field moves toward multi-agent systems, personalized assistants, and stateful copilots, we believe AFM offers a practical foundation for real-world memory management at scale.
Future work will explore learnable fidelity selection policies, hierarchical memory representations, and task-conditioned scoring functions, enabling AFM to evolve into a fully adaptive memory controller for next-generation language interfaces.
Acknowledgments
This work was conceived, implemented, and evaluated independently by the author.

### v3

@@
 Abstract.
-Large language models (LLMs) are increasingly deployed in multi-turn dialogue settings, but their behavior is still bottlenecked by fixed context windows and naïve memory strategies.
-Replaying the full conversation at every turn is simple but expensive, while static summarization or recency-only heuristics often erase safety-critical user details.
-We present Adaptive Focus Memory (AFM), a dynamic context manager that assigns each past message one of three fidelity levels—FULL, COMPRESSED, or PLACEHOLDER—based on semantic similarity to the current query, half-life recency weighting, and importance classification.
-AFM packs messages chronologically under a strict token budget, preferring high fidelity for the most relevant turns while aiming to preserve a cheap trace of the dialogue.
-In a safety-oriented benchmark involving a user with a severe peanut allergy planning a trip to Thailand, AFM retains the allergy across both short and medium-length conversations, matches the safety performance of naive replay, and cuts average token usage by 66% relative to a replay baseline.
-We release a modular Python implementation of AFM designed for OpenAI-compatible APIs and offline operation, enabling practitioners to reduce inference cost without sacrificing safety or factual continuity in the evaluated scenario.
+Large language models (LLMs) are increasingly deployed in multi-turn dialogue settings, yet their behavior remains bottlenecked by naive history-management strategies.
+Replaying the full conversation at every turn is simple but costly, while recency-based truncation or static summarization often causes early, high-impact user constraints to drift out of effective context.
+As a result, models may retain text without reliably applying it when it matters.
+We present Adaptive Focus Memory (AFM), a lightweight context management system that dynamically assigns each past message one of three fidelity levels—Full, Compressed, or Placeholder—based on semantic relevance, temporal decay, and importance classification.
+AFM packs messages chronologically under a fixed token budget, preserving critical constraints at high fidelity while allowing low-importance context to degrade gracefully.
+We evaluate AFM on two multi-turn dialogue benchmarks designed to stress long-horizon constraint preservation: a safety-critical travel scenario involving a user with a severe peanut allergy, and a policy-critical tax compliance scenario involving an illegal evasion request.
+Under strict grading that requires both explicit constraint recall and appropriately conditioned generation, AFM succeeds in 83.3 percent of allergy runs where all baseline strategies fail under strict grading, and preserves correct refusal behavior on the tax benchmark.
+These results demonstrate that effective dialogue memory requires more than retaining prior text.
+Selectively allocating fidelity across past messages enables reliable constraint preservation under bounded context growth, without modifying model weights or introducing external retrieval infrastructure.
+We release an open-source implementation of AFM compatible with OpenAI-style chat APIs to support reproducible research and practical deployment.
 1. Introduction
@@
   • We introduce a multi-fidelity replay mechanism that assigns each message to FULL, COMPRESSED, or PLACEHOLDER form and packs messages chronologically under a fixed token budget.
-  • We demonstrate that AFM retains critical user-specific facts—such as allergy declarations—more reliably than recency-only baselines, while matching the safety performance of naive replay at substantially lower cost.
+  • We demonstrate that AFM preserves safety-critical constraints more reliably than naive replay and recency-based heuristics under strict evaluation.
   • We release an open-source AFM implementationhttps://github.com/cruz209/AFMforLLM suitable for reproducible research and integration into production systems.
@@
 Several works propose architectures designed for local adaptation or long-context specialization.
-Meng et al.
-[7] introduce locally trainable memory layers for more efficient fine-tuning, while Geva et al.
+introduce locally trainable memory layers for more efficient fine-tuning, while Geva et al.
 [11] show that transformer feed-forward layers operate implicitly as key–value memory stores.
@@
 More recent surveys [8] highlight the difficulty of maintaining coherence when summarizing or retrieving historical dialogue, particularly when the user’s intent evolves across turns.
-Structured-memory architectures [13] attempt to improve retrieval fidelity, yet these systems require substantial infrastructure—vector databases, indexing, and latency-heavy retrieval pipelines.
+Structured-memory architectures attempt to improve retrieval fidelity, yet these systems require substantial infrastructure—vector databases, indexing, and latency-heavy retrieval pipelines.
 2.3. Summarization and Compression of Conversational History
@@
 It operates entirely at the prompt layer and is implemented as a simple Python library.
-  • Explicit compute minimization: Unlike long-context models or pure summarization strategies, AFM explicitly focuses on reducing token-level inference cost and reports these savings in standard units (tokens and dollars per million tokens).
 Taken together, AFM occupies a space between RAG, KV-extended transformers, and classical summarization—providing a memory framework whose primary objective is token-efficient factual continuity rather than scaling raw context length or storing external facts.
@@
 If no client is configured, all messages default to TRIVIAL and no classification calls are made.
-**AFM does not evaluate the accuracy of the importance classifier, its behavior is prompt-driven.**
+AFM does not evaluate the accuracy of the importance classifier, its behavior is prompt-driven.
 Final scoring rule.
@@
 A system preamble (if provided) is treated like any other message: it is included first, subject to the same budget accounting.
-3.6. System architecture and API.
+3.6. System architecture and API
 AFM is implemented as a modular Python package centered on a FocusManager class. The key methods are:
@@
 4.1. Objective and scope.
-Our experiments evaluate how effectively AFM heuristically balances three competing goals in multi-turn dialogue: (1) retaining user-specific factual constraints, (2) avoiding unsafe recommendations, and (3) minimizing token usage and latency.
-We focus on a safety-critical setup where the user has a severe peanut allergy and is planning a trip to Thailand.
+Our experiments evaluate whether AFM can reliably preserve high-importance constraints across multi-turn dialogue while controlling context growth.
+We focus on two safety-critical constraint types that appear frequently in production assistants but are not well covered by standard LLM benchmarks:
+  1. (1)
+    Personal safety constraints that must be remembered and applied later (e.g., a life-threatening food allergy).
+  2. (2)
+    Policy/legal constraints that must remain binding even after long benign context (e.g., refusing illegal tax-evasion assistance).
+Both settings target a relatively new problem space: context management as a first-class system component.
+Most existing evaluations either (i) assume the full conversation history is always available, or (ii) test the model in isolation without stressing the memory policy that a deployed system actually uses.
+In contrast, we evaluate AFM as a history-management strategy under long, mixed-topic conversations where the binding constraint appears early and is not restated verbatim.
+4.2. Why a custom benchmark.
+We found that off-the-shelf safety or instruction-following benchmarks do not cleanly isolate the problem AFM addresses: constraint drift under realistic, long dialogue with imperfect memory.
+In particular, we need:
+  • Constraint non-repetition: the constraint is stated once early, then later questions implicitly require it.
+  • Mixed benign distractors: intervening turns that are plausible in real usage but semantically distant.
+  • Strict pass semantics: it is not enough to merely mention the constraint; the response must treat it with appropriate seriousness.
+Accordingly, we use two scripted, multi-turn scenarios with automatic grading and full transcript logging.
 4.3. Task design.
-We construct a synthetic but realistic conversation where a user with a life-threatening peanut allergy asks about travel logistics, street food, and cultural experiences in Thailand.
-The allergy is stated early and then not repeated verbatim, while later questions implicitly require the model to remember and apply that constraint (e.g., “the street food sounds AWESOME I wanna have it all” and “What are the best street foods I should try”).
-Following the scripts in the benchmark repository, we consider two lengths of this scenario:
-  • Short (3 turns). The allergy and food-related queries are close together.
-  • Medium (9 turns). The allergy is mentioned early, followed by several intervening questions (e.g., about destinations, transport, and Muay Thai) before the final street-food query.
-In both cases, we evaluate whether each method remembers and correctly applies the peanut allergy when recommending food.
+We consider two benchmarks.
+Peanut allergy travel benchmark.
+A user planning a trip to Thailand states early: they have a severe peanut allergy.
+The conversation then proceeds through multiple benign topics (destinations, transport, cultural experiences), and culminates in a food recommendation query.
+A run passes if the final recommendation (i) recalls the allergy and (ii) conditions food guidance appropriately (e.g., avoiding peanut-heavy dishes, providing high-salience precautions).
+Tax compliance benchmark.
+A user provides an explicit constraint early: advice must comply with U.S.
+tax law and must not assist illegal evasion.
+After many benign tax questions, the user escalates to an explicit illegal request (hiding unreported freelance income).
+A run passes if the assistant clearly refuses and redirects toward legal, compliance-oriented alternatives.
 4.4. Compared methods.
-We compare AFM against three baselines that correspond directly to the provided benchmark scripts:
+We compare AFM against three baselines corresponding to our benchmark scripts:
   1. (1)
-    Default (stateless) chat.
-Only the current user turn and a fixed system prompt are sent to the model.
-No prior messages are included; there is no token budget enforcement.
-This is implemented in default.py.
+    Default (stateless) chat. Only the current user turn (plus a fixed system prompt) is sent.
   2. (2)
-    Naive truncated replay.
-All previous user and assistant messages are appended to the prompt, subject to a global token budget $B=800$.
-When the budget is exceeded, the oldest messages are dropped until the prompt fits.
-No compression is used; all retained messages are verbatim.
-This is implemented in naive_replay.py.
+    Naive truncated replay. Append prior turns until a global budget is hit, then drop oldest turns.
   3. (3)
-    Recency-based compression.
-The two most recent user–assistant turns are kept verbatim.
-All earlier messages are individually compressed to a fixed local budget (100 tokens per message) using LLMCompressor, and the compressed messages plus recent turns are concatenated without an additional global token cap.
-This is implemented in recency.py.
-AFM itself is instantiated via FocusManager with OpenAI embeddings and the LLMCompressor when an API key is present, matching the configuration used by the recency-based baseline.
+    Recency-based compression. Keep the most recent turns verbatim; compress older turns to a fixed local budget.
+AFM uses selective fidelity allocation via FocusManager: high-importance items are retained at Full fidelity, mid-importance items are compressed, and low-importance items are reduced to short stubs.
 4.5. Implementation details.
-All experiments use OpenAI’s gpt-4o-mini model as the backend for generation and, when applicable, for compression and importance classification. We use the following decoding temperatures:
-  • Default (stateless) baseline: temperature $0.0$ (deterministic).
-  • Naive truncated replay baseline: temperature $0.7$.
-  • Recency-based compression baseline: temperature $0.7$.
-  • AFM in our interactive demo: temperature $0.0$ for deterministic compression and context selection.
-Token usage in all baselines is measured with the shared TokenCounter configured for gpt-4o-mini.
-The truncated replay baseline explicitly uses a budget of 800 tokens when deciding which messages to retain; the recency-based baseline and the default baseline do not enforce a global budget and therefore can produce longer prompts.
-AFM’s hyperparameters in the OpenAI-backed experiments are:
-  • high-fidelity threshold $\tau_{\text{high}}=0.45$,
-  • medium-fidelity threshold $\tau_{\text{mid}}=0.25$,
-  • recency half-life $h=12$ turns,
-  • maximum stub length of 12 tokens,
-    These values are drawn directly from the FocusConfig in the reference code.
+All experiments use OpenAI’s gpt-4o-mini for generation.
+When enabled, the same model family is used for compression and importance classification.
+Token usage is measured with the shared tokenizer-backed TokenCounter.
+Latency is measured as wall-clock time from request submission to completion receipt.
 4.6. Evaluation metrics.
-    We assess each method along the following axes:
-    • Allergy retention. Whether the model explicitly recalls and applies the user’s peanut allergy when recommending food (binary pass / fail per conversation).
-    • Factual safety. Whether any recommended dishes clearly contain peanuts or common peanut-derived ingredients without appropriate warnings.
-    • Token usage. Total tokens consumed per conversation (system prompt, all context, and assistant responses), averaged across the short and medium runs.
-    • Latency. Wall-clock time per assistant turn, computed from prompt submission to receipt of the completion.
-    Token and latency metrics for the baselines are logged directly by the benchmark scripts. AFM token and latency measurements are collected using the same TokenCounter and timing conventions.
+Pass/fail semantics.
+For both benchmarks, we adopt a strict binary criterion:
+A run is a considered pass if and only if the model both (i) explicitly recalls the binding constraint and (ii) generates a response that treats it with appropriate seriousness.
+This intentionally separates memory failures (constraint not recalled) from generation failures (constraint recalled but handled weakly, e.g., unsafe dish suggestions with soft disclaimers).
+Reported metrics.
+For each method and benchmark (30 seeds each), we report:
+  • Pass rate (%).
+  • Token usage on the graded turn (mean $\pm$ std).
+  • Latency on the graded turn (mean $\pm$ std).
 4.7. Main results.
-    Table 1 reports representative results from a single short- and medium-length run per method under the configurations above.
-The numbers for the three baselines are computed directly from the provided benchmark logs; AFM’s numbers are obtained using the same measurement infrastructure in the AFM demo script.
-    Method Allergy Recall (S / M) Avg Tokens Latency (s) Safe?
-    Default (stateless) N / N 1,493 6.6 No
-    Naive truncated replay Y / Y 2,479 8.8 Yes
-    Recency compression Y / N 1,888 12.0 Potentially
-    AFM (ours) Y / Y 504 6.2 Yes
+Table 1 reports results on both benchmarks.
+Method Pass rate Tokens (mean $\pm$ std) Latency (s, mean $\pm$ std)
+Peanut allergy benchmark (Thailand travel)
+Default (stateless) 0/30 (0%) 194.0 $\pm$ 0.0 4.04 $\pm$ 0.77
+Naive truncated replay 0/30 (0%) 34.3 $\pm$ 46.5 11.49 $\pm$ 2.35
+Recency compression 0/30 (0%) 33.9 $\pm$ 45.5 11.33 $\pm$ 1.90
+AFM 25/30 (83.3%) 286.0 $\pm$ 0.0 21.24 $\pm$ 2.01
+Tax compliance benchmark (illegal evasion request)
+Default (stateless) 30/30 (100%) 44.0 $\pm$ 0.0 0.69 $\pm$ 0.17
+Naive truncated replay 30/30 (100%) 32.0 $\pm$ 0.0 0.78 $\pm$ 0.21
+Recency compression 30/30 (100%) 32.0 $\pm$ 0.0 0.75 $\pm$ 0.20
+AFM 30/30 (100%) 286.0 $\pm$ 0.0 20.84 $\pm$ 2.61
 Table 1.
-Benchmark results on short (S) and medium (M) conversations.
-Baseline numbers are calculated across the short and medium runs logged by the benchmark scripts; AFM is measured with the same tokenizer and timing setup.
-Reported metrics come from a single reference run and should be interpreted qualitatively, as AFM relies on external API calls (embedding, compression, importance classification) that make multi-run statistical evaluation costly and noisy due to API and network variance.
-A more comprehensive, multi-run evaluation is left for future work.
-Note that the recency-compression baseline does not enforce a global token budget and can exceed the 800-token cap that AFM and naive replay obey.
-This matches common production deployments of recency heuristics but means the baseline is not a strictly controlled comparison.
-Latency comparisons are approximate because decoding temperatures differ across baselines.
+Live benchmark results (30 seeds per benchmark).
+A pass requires both constraint recall and appropriate constraint-conditioned generation.
+AFM substantially improves constraint retention on the peanut allergy benchmark, where all baselines fail under strict grading.
+On the tax benchmark, all methods refuse the explicitly illegal request; AFM’s latency reflects additional system work (importance classification / context selection) beyond a single generation call.
+Key takeaways.
 Two patterns stand out.
-First, the stateless and recency-compressed baselines fail to reliably apply the allergy constraint in the medium-length scenario: they eventually drift to generic street-food recommendations that do not reflect the user’s stated constraint.
-Second, naive truncated replay maintains safety but at the cost of substantially larger prompts and higher latency.
-AFM achieves the same safety as naive replay on this benchmark while using considerably fewer tokens.
+(1) Personal-safety constraints are fragile under long dialogue.
+On the peanut allergy benchmark, all three baselines fail in every seed under strict grading.
+In the logged failures, baseline responses drift toward generic Thailand street-food recommendations and do not sustain the early allergy constraint strongly enough to be considered safe.
+In contrast, AFM passes 83.3% of runs by consistently re-surfacing the allergy declaration at high fidelity and conditioning food guidance on it.
+(2) Policy refusal is easier when the final user request is overt.
+On the tax benchmark, the final request is an explicit illegal evasion instruction.
+All methods refuse, which suggests that the limiting factor here is not long-range memory, but whether the system maintains a stable refusal policy across benign context.
+This benchmark therefore acts as a “sanity check” that AFM’s context selection does not drop or weaken the user-provided compliance constraint.
+(3) System-level trade-off: AFM adds overhead.
+In this live setup, AFM exhibits higher measured latency on the graded turn, consistent with additional pipeline steps (importance scoring, context selection, and optional compression).
+This is the expected systems trade: AFM spends computation to reduce constraint drift.
 4.8. Qualitative failure modes.
-    Inspecting the transcripts reveals characteristic failure modes:
-    • Default (stateless).
-Because the model only sees the current user message and a generic system prompt, it recommends canonical Thai street foods (including peanut-heavy dishes such as Pad Thai) without referencing the allergy, even when the allergy was stated earlier in the same conversation scenario under other methods.
-    • Recency-based compression.
-When the allergy happens to be within the last few turns, compressed context plus recent raw turns is often sufficient.
-Once the allergy falls outside the recent window, the model reverts to generic advice such as “if you have food allergies, ask vendors about ingredients,” without explicitly accounting for peanuts.
-    • Naive truncated replay.
-This baseline successfully preserves the allergy but often feeds the model long, repetitive prompts as the history grows.
-The resulting responses can be verbose and redundant, with little benefit over AFM in the benchmark scenario.
-    AFM, by contrast, assigns the allergy declaration a high importance score, securing it a Full slot even when older, and tends to keep the later food-related user turns at higher fidelity as well.
-This leads to explicit allergy-aware recommendations in both short and medium settings.
-    4.8. Ablation analysis (planned).
-    The current codebase exposes natural ablation dimensions—removing compression, disabling importance classification, altering the recency half-life, or swapping the LLM compressor for the heuristic one.
-However, we have not yet run a systematic ablation study.
-Formal ablations over these dimensions are left to future work and are not reported here.
-    4.9. Human evaluation.
-    We did not conduct a formal human evaluation with multiple annotators and pre-registered protocols.
-Instead, we relied on qualitative inspection of the logged transcripts across methods.
-In these spot checks, AFM’s responses were consistently more explicit about the peanut allergy in the final food recommendation turns, whereas the baselines often produced allergy-agnostic street-food lists once the allergy drifted out of their effective context window.
-    Because this inspection was informal and limited in scope, we do not report numerical human preference percentages. A more rigorous human evaluation is future work.
-    4.10. Compute cost analysis.
-    To relate token savings to monetary cost, we define a simple compute saving ratio (CSR).
-Let $T_{\text{base}}$ be the average number of tokens consumed by a baseline configuration and $T_{\text{AFM}}$ be the average tokens consumed by AFM on the same conversations.
-We define
-    $\mathrm{CSR}=\frac{T_{\text{base}}-T_{\text{AFM}}}{T_{\text{base}}}.$
-    Using the stateless default baseline as $T_{\text{base}}$ (average of short and medium runs), the empirical values in Table 1 yield $\mathrm{CSR}\approx 0.66$, i.e., AFM reduces token usage by roughly 65% relative to a default configuration that does not manage history at all.
-Relative to the replay-style baselines (naive truncated replay and recency compression), AFM achieves even larger fractional token reductions, on the order of 70–80% in this benchmark.
-    Assuming a linear cost model with price $C_{\text{perM}}$ USD per million input tokens—as is standard for commercial LLM APIs—the marginal cost reduction per million tokens is
-    $\Delta C=\mathrm{CSR}\times C_{\text{perM}}.$
-    For example, with $C_{\text{perM}}=\mathdollar 3$ and $\mathrm{CSR}\approx 0.66$, AFM would save roughly \$2 per million tokens relative to the default baseline.
-In high-throughput deployments, this cost difference compounds rapidly.
-    4.11. Summary.
-    On the evaluated allergy benchmark, AFM:
-    • preserves safety and allergy recall in both short and medium-length conversations;
-    • substantially reduces token usage compared to both a stateless default setup and replay-based baselines;
-    • avoids the extreme prompt bloat of naive replay while retaining key user-specific constraints more reliably than a simple recency-only compressor.
-    These results support the central claim of AFM: selectively allocating fidelity across past messages can yield strong safety and factual continuity and accuracy at significantly lower token cost than naive history management strategies.
+Inspection of transcripts reveals characteristic errors:
+  • Default (stateless): fails by construction on memory-dependent tasks; earlier constraints are never visible.
+  • Recency compression: succeeds only if the constraint remains within the recent raw window; once it ages out, the model falls back to generic disclaimers and unsafe or insufficiently cautious recommendations.
+  • Naive replay: is not a principled constraint-preservation mechanism: it can preserve early constraints when they remain in-budget, but becomes brittle as history grows and truncation deletes exactly the message that matters.
+AFM’s successes are consistent with its design: it assigns early constraint declarations high importance, promoting them to Full fidelity even when old, while allowing low-importance distractors to degrade to stubs.
+4.9. Ablation analysis.
+We ablate three AFM components on the peanut allergy benchmark:
+  1. (1)
+    No compression: disable compression of mid-importance items.
+  2. (2)
+    No placeholder/stubbing: disable stub-based degradation (retain higher-fidelity context more often).
+  3. (3)
+    No importance classification: disable the importance scorer (equivalent to removing the mechanism that locks critical constraints into Full fidelity).
+Table 2 reports pass rates and selected internal context statistics.
+Variant Pass rate Full count Compressed count Stub count
+AFM (full system) 25/30 (83.3%) 1.0 0.0 24.0
+No compression 24/30 (80.0%) 1.0 0.0 24.0
+No placeholders/stubs 21/30 (70.0%) 1.0 0.8 22.6
+No importance classification 0/30 (0%) 0.0 0.0 26.0
+Table 2.
+AFM ablations on the peanut allergy benchmark (30 seeds).
+Disabling importance classification collapses performance to 0%, indicating that which messages receive Full fidelity is the dominant driver of constraint preservation.
+Other ablations reduce pass rate primarily via generation-quality failures: the allergy is often mentioned but treated too softly.
 5. Discussion
-5.1. When Context Is Not Comprehension
-One of the core findings in our experiments is that merely including prior messages in an LLM’s prompt does not guarantee effective memory.
-The naive replay strategy—common in many production chat systems—preserves all text but still fails to consistently apply user-specific facts when they matter.
-In our allergy benchmark, stateless and recency-compressed configurations ignored life-critical details (e.g., a peanut allergy) just one or two turns after they appeared.
-This reinforces a central problem in long-context inference: information salience degrades with token distance.
-LLMs, even with extended context windows, exhibit a recency bias and a tendency to overweight immediately proximal text.
-This creates an illusion of memory where none exists and highlights the need for systems that explicitly curate and scaffold model attention over time.
-5.2. AFM’s Role in the Context Stack
-AFM intervenes at the level of prompt construction—neither modifying the model weights nor introducing complex retrieval infrastructure.
-This makes it suitable for developers seeking efficient memory management without sacrificing simplicity, interpretability, or control.
-Whereas RAG pipelines introduce database latency and risk discourse in coherence due to chunk fragmentation, AFM operates inline, restructuring the prompt without external calls.
-Compared to static summarization buffers (e.g., fixed-length history windows or periodic compress-and-truncate), AFM dynamically rescores every message per query, allowing real-time adaptation to user intent.
-In this sense, AFM functions as a soft attention mechanism over dialogue history—prioritizing based on both semantic relevance and temporal context, and deciding how “heavy” each memory needs to be: full, compressed, or placeholder.
-AFM clamps cosine similarity at zero, discarding negative similarity information.
-5.3. Efficiency Without Blind Compression
-One of AFM’s key contributions is demonstrating that memory compression need not be all-or-nothing.
-Instead of summarizing the entire past into a single global buffer (which risks flattening important details), AFM preserves structure.
-Every message is individually scored and passed through a fidelity filter, enabling nuanced context packing.
-Empirically, this approach achieves strong token efficiency—reducing prompt size by over 60% on average—without compromising on safety-critical memory.
-AFM outperforms both naive replay and recency-only baselines on safety recall and token efficiency, and achieves competitive latency in our evaluated scenario.
-5.4. Failure Modes and Challenges
-AFM, while promising, is not immune to edge cases. Several challenges merit deeper scrutiny:
-  • Semantic drift from LLM compression.
-Abstractive summarization is inherently lossy.
-While effective for reducing tokens, poor LLM summaries may misrepresent prior user intent or omit subtle qualifiers—especially for long or multi-part messages.
-Hallucinated details, although rare in our benchmark, remain a risk under low-temperature decoding.
-  • Overrepresentation of low-information turns.
-Messages with high embedding similarity but low informational value (e.g., “sure,” “got it” ) may receive elevated scores unless filtered via length or entropy thresholds.
-Future versions could incorporate dialogue-act classification to suppress such filler turns.
-  • Interaction effects across messages.
-AFM scores each message independently, ignoring interactions like pronoun chains or deferred clarifications.
-This may result in dropped context for referential utterances unless earlier turns are retained.
-A graph-based salience model might help capture such dependencies.
-  • Static hyperparameters.
-Current AFM settings use fixed half-life and importance multipliers.
-Adaptive weighting (e.g., task-conditioned recency schedules) could better balance short-term versus long-term salience.
-5.5. Toward Learned or Reinforced Memory Policies
-AFM’s current decision logic is greedy and rule-based. However, as context management becomes more central to multi-agent systems, we envision extensions that learn optimal memory policies:
-  • Reinforcement-learned packing. Agents could be rewarded for efficient context usage that preserves factual recall or reduces hallucination rate.
-  • Memory trees. Instead of a flat log, context could be stored in a hierarchical structure with collapsible nodes and rewritable summaries.
-  • Personalization-aware retention.
-Messages may be scored differently based on user identity or long-term intent profiles.
-This would require coupling AFM with identity tracking or goal-prediction modules.
-5.6. Broader Implications
-As LLMs are increasingly embedded in tool-using agents, IDE copilots, or multi-turn reasoning chains, the need for cost-aware, interpretable, and controllable memory systems is growing.
-AFM’s design makes it especially attractive in the following real-world scenarios:
-  • Conversational agents with safety constraints. Applications like healthcare chatbots or legal assistants must retain critical facts across turns without compromising on latency or privacy.
-  • Agentic workflows with recurrent state. Planning agents or goal-oriented bots benefit from a memory that preserves task-relevant state across tool invocations.
-  • Model distillation and compression pipelines. AFM can serve as a teacher signal in training smaller models to simulate long-context behavior without full replay.
-In all these contexts, the need is the same: not just to remember more, but to remember better. AFM’s selective fidelity framework takes a concrete step toward that goal.
+5.1. Context Availability Does Not Imply Constraint Use
+Across all non-AFM baselines, we observe a consistent failure to apply user-specified constraints at the point where they become decision-critical.
+In the live peanut-allergy benchmark, Default-Live, Naive Replay, and Recency-weighted strategies all achieve a 0% pass rate across seeds under the safety judge.
+These failures arise from distinct, but related, limitations in how context is handled.
+In Default-Live, each turn is evaluated independently, and the final query is issued without access to prior dialogue.
+As a result, the allergy constraint is not present in the model’s context window at the final turn, and cannot influence the response.
+In contrast, Naive Replay and Recency-weighted approaches frequently retain the allergy disclosure within the context window at the final turn.
+Under naive replay, the constraint may appear verbatim when it survives truncation, while under recency-weighted compression it is typically present in a condensed form.
+Despite this retention, neither approach succeeds in enforcing the constraint: across all seeds, the model fails to consistently acknowledge the allergy, provide appropriate safety guidance, or avoid unsafe recommendations.
+This distinction highlights a critical point: constraint failure is not solely a consequence of context truncation.
+Even when safety-critical information remains available within the prompt, it often lacks sufficient salience to affect model behavior.
+The result is an illusion of memory, in which information is retained textually but remains operationally inert.
+5.2. AFM and Constraint Reactivation
+Adaptive Focus Memory (AFM) addresses this failure mode by explicitly managing how prior information is represented as interaction unfolds.
+Rather than treating dialogue history as a flat buffer, AFM dynamically reallocates prompt budget across memory items based on inferred importance, assigning each item a representation of appropriate fidelity.
+Empirically, this difference is decisive.
+In the same live peanut-allergy benchmark where all non-AFM baselines achieve 0% pass rates, AFM achieves a substantially higher rate of constraint adherence across seeds.
+Unlike baseline methods, AFM consistently preserves the allergy constraint in a high-salience form at the final turn, enabling the model to acknowledge the allergy, prioritize safety, and avoid hazardous recommendations.
+Crucially, AFM does not require a larger context window, modify model weights, or introduce external retrieval.
+All improvements arise from restructuring the prompt itself.
+By ensuring that critical constraints remain both present and prominent when relevant, AFM enables what prior methods fail to provide: reactivation of earlier information at decision time.
+5.3. Ablation Analysis: The Role of Selective Fidelity
+Ablation results further indicate that AFM’s gains cannot be attributed to any single heuristic.
+Removing placeholder retention, adaptive compression, or importance-based scoring each leads to marked degradation in performance.
+In particular, ablations that retain compressed summaries of constraints but remove explicit placeholder markers revert toward baseline behavior.
+This suggests that compressed representations alone are insufficient unless the system explicitly signals their continued relevance.
+Conversely, disabling compression while preserving importance scoring increases token usage without recovering reliability.
+These findings support the view that AFM’s effectiveness derives from selective fidelity: allocating limited prompt space unevenly, so that safety-critical facts remain sharply represented while less relevant history is progressively abstracted.
+5.4. Implications for Long-Horizon Interaction
+Taken together, these results suggest that failures in long-horizon LLM behavior should not be understood primarily as failures of context length.
+Instead, they reflect a mismatch between how context is stored and how it is later used.
+Baseline strategies demonstrate two failure modes: absence of relevant information (Default-Live) and presence without influence (Naive Replay and Recency-weighted).
+AFM resolves both by treating context as a dynamically managed resource rather than a static transcript.
+More broadly, this work suggests that improving long-horizon reliability will require systems that explicitly manage salience, abstraction, and persistence of constraints over time.
+AFM represents one such approach, and its performance across safety-critical scenarios indicates that explicit context management is a necessary component of robust interactive language systems.
 6. Conclusion
@@
 Through a combination of embedding-based scoring, half-life recency weighting, and hybrid compression techniques, AFM preserves the key factual constraint in our evaluated scenario while aggressively reducing prompt size.
-Our experiments show that AFM outperforms baseline approaches on key metrics of factual retention, safety alignment (AFM matches naive replay on safety alignment in our evaluated scenario at a much lower token cost.), and token efficiency in a safety-critical benchmark.
+Our experiments show that AFM outperforms baseline approaches on key metrics of factual retention, safety alignment in a safety-critical benchmark.
 Unlike retrieval-augmented or static summarization methods, AFM operates inline, requires no external infrastructure, and adapts fluidly to changing user intent.